OpenStudy (anonymous):

a cone with height h is inscribed a larger cone with height 12 so that its vertex is at the center of the base of the larger cone. find height of inner cone that maximizes its volume

OpenStudy (aum):



OpenStudy (aum):



OpenStudy (aum):

Use similar triangles to find r in terms R

OpenStudy (aum):

OCD and OAB are similar triangles. OC / OA = r / R (12-h) / 12 = r / R r = (12-h)/12 * R Volume of inner cone V = 1/3 * pi * r^2 * h r = (12-h)/12 * R Put this in V above and maximize V wrt h. (Find dV/dh and equate it to zero and solve for h).